At sea level, the weight of the atmosphere exerts a pressure of 14.7 pounds per square inch, commonly referred to as 1 atmosphere of pressure. As an object descends in water, pressure P and depth d are linearly related. In salt water, the pressure at a depth of 33 ft is 2 atms, or 29.4 pounds. per square inch.
1) Choose the correct linear model below.
__
a) P = 0.445d + 14.7
b) P = 2.245d - 33
__
c) P = 0.445d + 29.4
__
d) P = 0.03d + 1

2) Choose the correct interpretation of the slope of the model below
a) It is the rate of change of pressure with respect to depth in units of lb/inē per foot
b) It is the rate of change of depth with respect to pressure in units of feet per lb/inē


3) What is the pressure at a depth of 90 ft?
______ pounds per square inch (Round to the nearest integer as needed.)

4) What is the depth at which the pressure is 6 atms?
____ feet (simplify your answer.)

Answer by ankor@dixie-net.com(22740)

At sea level, the weight of the atmosphere exerts a pressure of 14.7 pounds per square inch, commonly referred to as 1 atmosphere of pressure. As an object descends in water, pressure P and depth d are linearly related. In salt water, the pressure at a depth of 33 ft is 2 atms, or 29.4 pounds. per square inch.
1) Choose the correct linear model below.
__
a) P = 0.445d + 14.7
Found by 14.7/33 = .445, plus the pressure of the air
:
2) Choose the correct interpretation of the slope of the model below
a) It is the rate of change of pressure with respect to depth in units of lb/inē per foot
:
3) What is the pressure at a depth of 90 ft?
______ pounds per square inch (Round to the nearest integer as needed.)
P = .445(90) + 14.5
P = 55 psi
4) What is the depth at which the pressure is 6 atms?
____ feet (simplify your answer.)
Find the pressure at 6 atms: 6 * 14.7 = 88.2
.445d + 14.7 = 88.2
.445d = 88.2 - 14.7
.445d = 73.5
d = 73.5/.445
d = 165 ft is the depth at 6 atms

0 solutions